# Bandits with Preference Feedback: A Stackelberg Game Perspective

### v2

@@ D.1 Yelp Experiment
We filter the Yelp data [^3] Source: https://www.yelp.com/dataset for restaurants in Philadelphia, USA with at least $500$ reviews and users who reviewed at least $90$ restaurants. The final dataset includes $275$ restaurants, $20$ users, and a total of $2563$ reviews. We define the action space ${\mathcal{X}}$ by assigning to each restaurant their respective $32$-dimensional embedding of their reviews, i.e., ${\mathcal{X}}\subseteq{\mathbb{R}}^{32}$. For each restaurant, we concatenate all reviews in the filtered dataset and we use the text-embedding-3-large OpenAI embedding model  [^4] Source: https://platform.openai.com/docs/guides/embeddings to retrieve the embeddings. The Yelp dataset provides utility values for users in the form of ratings on the scale of $1$ to $5$, however, not all users rated every restaurant. We use collaborative filtering to estimate the missing reviews [Schafer et al., 2007]. For each user, these values are used as the utilities for restaurants during the simulation. Experiments are conducted separately for each user, therefore, utility values are not aggregated but the action space is identical for each experiment.
Note that we do not assume any explicit functional form of the utility functions $f$ that we calibrate to this data. Instead, the actions space ${\mathcal{X}}$ and utility values are derived separately from the dataset. Regardless, the results presented in Section 6.3 show that our kernelized method achieves good performance on this task.
@@ Appendix E Additional Experiments
In this section, we provide Table 2 that details our logistic bandit benchmark, complementing the results in Section 6.1. Figure 4 and Figure 9 show the logistic and dueling regret of additional test functions, complementing the results of Table 1.
@@
Figure 4: Regret with Branin utility function with logistic and preference feedback.
Figure 9: Top to bottom Regret for Eggholder, Hölder, Matyas Michalewicz, Rosenbrock functions, with logistic and preference feedback.

### v3

@@ D.1 Yelp Experiment
We filter the Yelp data [^3] Source: https://www.yelp.com/dataset for restaurants in Philadelphia, USA with at least $500$ reviews and users who reviewed at least $90$ restaurants. The final dataset includes $275$ restaurants, $20$ users, and a total of $2563$ reviews. We define the action space $\mathcal{X}$} by assigning to each restaurant their respective $32$-dimensional embedding of their reviews, i.e., ${\mathcal{X}}\subseteq{\mathbb{R}}^{32}$. For each restaurant, we concatenate all reviews in the filtered dataset and we use the text-embedding-3-large OpenAI embedding model  [^4] Source: https://platform.openai.com/docs/guides/embeddings to retrieve the embeddings. The Yelp dataset provides utility values for users in the form of ratings on the scale of $1$ to $5$, however, not all users rated every restaurant. We use collaborative filtering to estimate the missing reviews [schafer2007collaborative]. For each user, these values are used as the utilities for restaurants during the simulation. Experiments are conducted separately for each user, therefore, utility values are not aggregated but the action space is identical for each experiment.
Note that we do not assume any explicit functional form of the utility functions $f$ that we calibrate to this data. Instead, the actions space $\mathcal{X}$} and utility values are derived separately from the dataset. Regardless, the results presented in Section˜6.3 show that our kernelized method achieves good performance on this task.
@@ Appendix E Additional Experiments
In this section, we provide Table˜2 that details our logistic bandit benchmark, complementing the results in Section˜6.1. Figure˜4 and Figure˜5 show the logistic and dueling regret of additional test functions, complementing the results of Table˜1.
@@
Figure 4: Regret with Branin utility function with logistic (left) and preference (right) feedback.
Figure 5: Top to bottom Regret for Eggholder, Hölder, Matyas Michalewicz, Rosenbrock functions, with logistic (left) and preference (right) feedback.